# CLEAR: Towards Contextual LLM-Empowered Privacy Policy Analysis and Risk Generation for Large Language Model Applications

### v2

@@ Abstract.
The rise of end-user applications powered by large language models (LLMs), including both conversational interfaces and add-ons to existing graphical user interfaces (GUIs), introduces new privacy challenges. However, many users remain unaware of the risks. This paper explores methods to increase user awareness of privacy risks associated with LLMs in end-user applications. We conducted five co-design workshops to uncover user privacy concerns and their demand for contextual privacy information within LLMs. Based on these insights, we developed CLEAR (Contextual LLM-Empowered Privacy Policy Analysis and Risk Generation), a just-in-time contextual assistant designed to help users identify sensitive information, summarize relevant privacy policies, and highlight potential risks when sharing information with LLMs. We evaluated the usability and usefulness of CLEAR across in two example domains: ChatGPT and the Gemini plugin in Gmail. Our findings demonstrated that CLEAR is easy to use and improves user understanding of data practices and privacy risks. We also discussed LLM’s duality in posing and mitigating privacy risks, offering design and policy implications.
@@ 1. Introduction
In this paper, we focus exclusively on scenarios where users directly interact with LLMs, excluding those scenarios in which LLMs are used behind the scenes to analyze user data (Ma et al., 2023; Dai et al., 2023) or substitute traditional machine learning models in back-end processes (Tan et al., 2023). Our analysis considers two most popular user interactions with LLMs – when users directly interact with LLMs through conversational interfaces (e.g., ChatGPT) and when users interact with LLM-based add-ons in existing graphical user interface (GUI) applications (e.g., Gemini plugin in Gmail, Copilot in Microsoft Office).
Despite the emergence of these privacy risks, most users either remain unaware or misunderstand these issues (Zhang et al., 2024). Users tend to trust LLMs like ChatGPT due to their human-like interactions and high utility, making them more likely to disclose personal information (Zhang et al., 2024). They often prioritize concerns about intellectual property leaks over personal privacy risks (Zhang et al., 2024). As a result, users frequently share private (e.g., emails) or semi-private (e.g., Facebook group posts) information with LLMs (Li et al., 2024). For instance, Zhang et al. identified numerous pieces of private data—including emails, phone numbers, and locations—in the publicly available ShareGPT52K dataset, which contains 50,496 ChatGPT chat histories (Zhang et al., 2024). Similarly, a report from Cyberhaven revealed that 8.6% of 1.6 million workers pasted confidential data into ChatGPT between its launch and June 2023 (Coles, 2023).
@@
However, these efforts overlook a critical aspect—enhancing users’ awareness of how their data is being used by LLMs and helping them understand and reflect on the reasons and risks associated with such usage. This is challenging because, firstly, users tend to prioritize the immediate utility and convenience offered by these systems over the long-term consequences of sharing personal information. The immediate benefits from interacting with LLMs frequently eclipse more abstract concerns such as privacy. This discrepancy leads to an overestimation of the model’s safety and an underestimation of its potential to inadvertently expose private data (Zhang et al., 2024). Additionally, many users lack sufficient privacy literacy, which hampers their ability to comprehend the technical details of data usage, model training, and the risks associated with data memorization (Gumusel et al., 2024).
@@
We evaluated its usability and usefulness in two distinct use cases. The first involved a browser plugin for ChatGPT, exemplifying standalone conversational interfaces like Claude and Gemini. In this study, 13 participants engaged in scenarios requiring them to input sensitive data into ChatGPT. The second use case utilized the Gemini addon within Gmail, exemplifying “add-on” LLM tools in existing GUIs, similar to Microsoft Copilot for Office applications. Here, 15 participants used the Gemini plugin in Gmail to reply to or summarize emails containing sensitive information. The results indicated that CLEAR significantly enhanced the user experience and increased awareness among participants regarding data practices and privacy risks associated with LLMs. Encouraged by their newfound knowledge of the potential risks and the respective privacy policies, most participants would chose to either delete the sensitive information or substitute it with synthesized fake data, leading to improved privacy management behaviors.
@@
  • We outline design and policy recommendations that aim to enhance privacy awareness, knowledge, and informed decision-making in user interactions with LLMs.
@@ 2.1. Privacy Risks of LLMs
Privacy risks of LLMs is a critical problem that hampers their adoption. Several privacy incidents have occurred recently. For example, in March 2023, a system bug of ChatGPT leaked users’ conversation history and payment information (Kshetri, 2023). Beyond conventional data breaches, three new types of new privacy challenges associated with LLMs have emerged: memorization risks, inference risks, and disclosure risks.
Memorization risks. LLMs are trained from vast amounts of data, including user-provided data through interactions. Therefore, these models can memorize and unintentionally generate users’ sensitive or private information from its training set (Inan et al., 2021; Lukas et al., 2023). Even with stricter prompt review mechanisms by LLMs, previous studies still indicated that malicious parties could execute a training data extraction attack and obtain or infer specific sensitive information through the fine-tuning of jailbreaking prompts (Li et al., 2023) or simulated guess game (Staab et al., 2023).
Inference risks. Inference risks refer to the ability of LLMs to automatically deduce a wide range of personal attributes about an author from seemingly innocuous text, due to their advanced inference capabilities (Li et al., 2024). For instance, Stabb et al. (Staab et al., 2023) shows that a phrase like “I always get stuck there waiting for a hook turn” can enable malicious actors to infer that the individual is in Melbourne, as the term “hook turn” is specific to that location. Although text anonymizing tools provide some level of protection, LLMs can often detect more subtle linguistic cues and contextual details that reveal personal information.
Disclosure risks. LLMs can now communicate with users in a human-human communication style through natural language. This anthropomorphic experience may inflate users’ trust and expectation (Kim and Sundar, 2012; Złotowski et al., 2015), which lead to greater disclosure of private information to LLMs. For example, Zhang et al. (Zhang et al., 2024) found that human-like interactions encourage users to unintentionally share sensitive and personally identifiable information with LLM-based conversational agents. They also noted that users believed they had to “pay the price” of privacy to get benefits from LLMs. This belief, coupled with manipulative interface designs (i.e., dark patterns (Lu et al., 2024)) that seek to persuade users into permit access to their privacy information, intensifies the risks of disclosure.
The risks mentioned above all originate from users providing sensitive data, either directly or indirectly, to LLMs. To address this issue, we aim to develop a tool that enhances users’ understanding of relevant privacy policies and the potential risks associated with their data before they share it.
@@ 2.2. Existing Methods for Protecting User Privacy when using LLM-enabled Applications
To protect user privacy when using LLM-enabled end-user applications, existing methods are mainly divided into two categories: model-side and user-side.
Model-side solutions can take effect in three stages (Smith et al., 2023): pre-processing, training, and post-training. In the pre-processing phase, data sanitization is a widely used technique to remove sensitive information from training data (Lison et al., 2021; Vakili et al., 2022). Given the large volume of training data, it is only practical to use automated methods (e.g., pattern-based parsing) to identify and remove specific sensitive information. However, defining private information is challenging due to its context-dependent nature and varied formats, making it difficult to guarantee complete sanitization (Brown et al., 2020). In the training stage, differential privacy (Dwork et al., 2016) is a commonly used technique for protecting user privacy. It works by adding random noise to the training data, making it difficult to identify any specific individual’s information while still allowing useful insights to be extracted. Although applying a differential privacy algorithm is effective in preventing memorization (Downey et al., 2022; Ponomareva et al., 2023), it also causes degradation of model performance and increased computation and memory usage (Ishihara, 2023). In the post-training stage, filtering and knowledge unlearning are two primary methods for handling sensitive information. Filtering involves identifying and removing outputs that contain sensitive information before they reach users. For example, Perez et al. (Perez et al., 2022) proposed a method to automatically generate test cases and identify potentially dangerous outputs from large language models (LLMs). However, identifying items to be filtered incurs a cost, and ensuring output diversity remains challenging (Ishihara, 2023). Knowledge unlearning (Jang et al., 2022) refers to methods that force models to forget specific knowledge within LLMs. The success of this method is highly dependent on the specific target data required for the model to be unlearn (Smith et al., 2023) and therefore the performance can fluctuate in different use cases.
User-side solutions take into account user engagement and interaction. For example, Kim et al. (Kim et al., 2024) proposed ProPILE, a privacy probing tool for investigating personally identifiable information (PII) leakage in LLMs. ProPILE allows users to input a subset of their PII into LLMs and test whether LLMs can predict or reveal the remaining PII of the user. Although it can provide users with information regarding possibility of different PII leaks in LLMs, using such a tool often disconnects from the actual context in which users use LLMs. To resolve this issue, Dou et al. (Dou et al., 2023) introduced a self-disclosure abstraction model that can paraphrase user input sensitive information into more generic terms while preserving their utility. For example, an input “I live in New Mexico” could be abstracted to “I live in the Southwest of the United States,” reducing the specificity of the information while maintaining its general meaning. Similar to digital nudges (Ioannou et al., 2021), this method can reduce the chance of data disclosure in LLM, but users still remain unaware of who might access the sensitive information, why it is being accessed, and the potential privacy risks.
Compared to previous work, we focus on enhancing end-users’ privacy awareness when using LLM-enabled applications by providing them with real-time contextual information. By informing users with privacy policy snippets and potential privacy policies related to the specific sensitive information they input, users can gain a more comprehensive understanding of the potential consequences of sending sensitive information to LLMs, enabling them to make more informed decisions.
@@ 2.3. Contextual Privacy Policy
Contextual privacy policy (CPP) deconstructs length privacy notices into shorter and digestible segments, displaying them only in the relevant context of use (Feth, 2017; Windl et al., 2022). Bergmann (Bergmann, 2008) showed that presenting only relevant information from privacy policies can significantly increase users’ privacy awareness. This also aligns with Nissenbaum’s theory of contextual integrity (Nissenbaum, 2004), which emphasizes that context is critical to determine whether a specific action is a violation of privacy.
Previous work has investigated CPP in both web and mobile systems. In the general web domain, Ortlof et al. (Ortloff et al., 2020) develop a concept showcase of CPP for seven different websites and collect the design implications of CPP. Building on it, Windl et al. (Windl et al., 2022) presented PrivacyInjector, a more scalable system that can automatically generate and display CPPs. In the mobile domain, Pan et al. (Pan et al., 2024) proposed SeePrivacy, which integrates vision-based GUI understanding with privacy policy analysis to automatically generate CPPs for mobile applications.
Compared with previous work, we focus on the new unique challenges of applying CPP in the area of LLM-enabled end-user applications. We not only generate CPP related to users’ input sensitive information but also leverage LLM to envision potential privacy risks associated with LLMs, enabling users to fully understand the potential implications and consequences of sharing the sensitive information that they are about to send out.
@@ 3. Co-Design Workshops
To understand user needs and explore the design space of tools to inform participants of privacy risks in LLM-enabled end-user applications, we conducted five sessions of co-design workshops [^1] The protocol of the workshop has been reviewed and approved by the IRB at our institution., including three in-person sessions and two online sessions.
@@ 3.1. Participants
“As noted above, we may use Content you provide us to improve our Services, for example to train the models that power ChatGPT.”
@@ 3.2. Workshop Procedure
Each co-design workshop began with an introductory session, including ice-breaker activities to foster a collaborative atmosphere. Participants then engaged in a discussion on privacy by reading three news articles on incidents relevant to LLMs (e.g., an article entitled “ChatGPT maker OpenAI faces a lawsuit over how it used people’s data [^2] https://www.washingtonpost.com/technology/2023/06/28/openai-chatgpt-lawsuit-class-action/ from the Washington Post)”. They identified potential privacy risks and discussed strategies to mitigate them. This initial activity aimed to build a foundational understanding of privacy challenges associated with LLMs.
@@
Following this, participants engaged in an ideation session where they shared their concerns and opportunities regarding privacy protections. This session led into three focused activities where participants crafted privacy policy elements, assessed privacy risks, and explored potential privacy incidents through sketching, writing, and group discussions. Table 2 shows the example materials used in the co-design workshop.
@@
In the first activity, participants designed the content and presentation of retrieved privacy policy snippets that were relevant to sensitive information in example scenarios, based on an example prompt input into the ChatGPT system. Participants could either write descriptions of the interface features or create sketches of the interfaces. The purpose of this ideation activity was to gather user-specific requirements for contextual privacy policies in the context of LLM-enabled applications.
@@
Building on the first activity, the second involved participants examining not only the initial prompt and the resulting privacy policy snippets but also potential privacy risks regarding sharing sensitive information with the LLM. The privacy risks were selected based on Lee et al.’s taxonomy of AI privacy risks (Lee et al., 2024) and were presented through detailed descriptions and visual aids, including examples of how personal information entered by users could be shared, used for advertising, or stored by service providers. The participants discussed and designed how to organize and present this information to better raise users’ awareness and support informed actions during a user’s interaction with the LLM-enabled applications. The activity specifically focused on the realistic context of use and how such information can fit into the user’s overall task that they are trying to achieve through the use of LLM-enabled applications. The goal was to understand whether highlighting LLM-related privacy risks could enhance users’ privacy awareness, and if so, how.
@@
The final activity focused on the presentation of specific privacy incidents, building on the risks identified earlier. Participants used the same initial prompt and related real-world incidents to design how these could be presented to heighten privacy awareness. The objective was to understand the impact of explicit privacy incident examples on participants’ awareness and to determine effective methods for communicating these incidents within the context of LLM application use.

### v3

@@ Abstract.
The rise of end-user applications powered by large language models (LLMs), including both conversational interfaces and add-ons to existing graphical user interfaces (GUIs), introduces new privacy challenges. However, many users remain unaware of the risks. This paper explores methods to increase user awareness of privacy risks associated with LLMs in end-user applications. We conducted five co-design workshops to uncover user privacy concerns and their demand for contextual privacy information within LLMs. Based on these insights, we developed CLEAR (Contextual LLM-Empowered Privacy Policy Analysis and Risk Generation), a just-in-time contextual assistant designed to help users identify sensitive information, summarize relevant privacy policies, and highlight potential risks when sharing information with LLMs. We evaluated the usability and usefulness of CLEAR across two example domains: ChatGPT and the Gemini plugin in Gmail. Our findings demonstrated that CLEAR is easy to use and improves users’ understanding of data practices and privacy risks. We also discussed LLM’s duality in posing and mitigating privacy risks, offering design and policy implications.
large language model, privacy awareness, privacy intervention, privacy literacy
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 30th International Conference on Intelligent User Interfaces; March 24–27, 2025; Cagliari, Italybooktitle: Proceedings of the 30th International Conference on Intelligent User Interfaces (IUI ’25), March 24–27, 2025, Cagliari, Italydoi: 10.1145/3708359.3712156ccs: Security and privacy Privacy protectionsccs: Human-centered computing Interactive systems and tools
@@ 1. Introduction
In this paper, we focus exclusively on scenarios where users directly interact with LLMs, excluding those scenarios in which LLMs are used behind the scenes to analyze user data (Ma et al., 2023; Dai et al., 2023) or substitute traditional machine learning models in back-end processes (Tan et al., 2023). Our analysis considers two most popular user interactions with LLMs—when users directly interact with LLMs through conversational interfaces (e.g., ChatGPT) and when users interact with LLM-based add-ons in existing graphical user interface (GUI) applications (e.g., Gemini plugin in Gmail, Copilot in Microsoft Office).
Despite the emergence of these privacy risks, most users either remain unaware or misunderstand these issues (Zhang et al., 2024; Wang et al., 2025). Users tend to trust LLMs like ChatGPT due to their human-like interactions and high utility, making them more likely to disclose personal information (Zhang et al., 2024). They often prioritize concerns about intellectual property leaks over personal privacy risks (Zhang et al., 2024). As a result, users frequently share private (e.g., emails) or semi-private (e.g., Facebook group posts) information with LLMs (Li et al., 2024a). For instance, Zhang et al. identified numerous pieces of private data—including emails, phone numbers, and locations—in the publicly available ShareGPT52K dataset, which contains 50,496 ChatGPT chat histories (Zhang et al., 2024). Similarly, a report from Cyberhaven revealed that 8.6% of 1.6 million workers pasted confidential data into ChatGPT between its launch and June 2023 (Coles, 2023).
@@
However, these efforts overlook a critical aspect—enhancing users’ awareness of how their data is being used by LLMs and helping them understand and reflect on the reasons and risks associated with such usage. This is challenging because, firstly, users tend to prioritize the immediate utility and convenience offered by these systems over the long-term consequences of sharing personal information. The immediate benefits from interacting with LLMs frequently eclipse more abstract concerns such as privacy. This discrepancy leads to an overestimation of the model’s safety and an underestimation of its potential to inadvertently expose private data (Zhang et al., 2024). Additionally, many users lack sufficient privacy literacy, which hampers their ability to comprehend the technical details of data usage (Wang et al., 2025), model training, and the risks associated with data memorization (Gumusel et al., 2024).
@@
We evaluated its usability and usefulness in two distinct use cases. The first involved a browser plugin for ChatGPT, exemplifying standalone conversational interfaces like Claude and Gemini. In this study, 13 participants engaged in scenarios requiring them to input sensitive data into ChatGPT. The second use case utilized the Gemini addon within Gmail, exemplifying “add-on” LLM tools in existing GUIs, similar to Microsoft Copilot for Office applications. Here, 15 participants used the Gemini plugin in Gmail to reply to or summarize emails containing sensitive information. The results indicated that CLEAR significantly improved the user experience and increased the awareness among participants regarding data practices and privacy risks associated with LLMs. Encouraged by their newfound knowledge of the potential risks and the respective privacy policies, most participants would choose to either delete the sensitive information or substitute it with synthesized fake data, leading to improved privacy management behaviors.
@@
  • We outline design and policy recommendations to enhance privacy awareness, knowledge, and informed decision-making in user interactions with LLMs.
@@ 2.1. Privacy Risks of LLMs
Privacy risks of LLMs are a critical problem that hampers their adoption. Notable incidents, such as the March 2023 ChatGPT bug that leaked user conversation histories and payment information (Kshetri, 2023), highlight these concerns. Beyond conventional data breaches, three new types of privacy challenges associated with LLMs have emerged: memorization risks, inference risks, and disclosure risks.
Memorization risks. LLMs are trained from vast amounts of data, including user-provided data through interactions. Therefore, these models can memorize and unintentionally reproduce sensitive information from their training data (Inan et al., 2021; Lukas et al., 2023). Even with strict prompt reviews, attackers can exploit models through training data extraction attacks or jailbreaking prompts (Li et al., 2023; Staab et al., 2023).
Inference risks. Inference risks refer to LLMs’ ability to automatically deduce a wide range of personal attributes about individuals from seemingly innocuous text, due to their advanced inference capabilities (Li et al., 2024a). For example, phrases like “waiting for a hook turn” can reveal a user is in Melbourne, as “hook turn” is location-specific (Staab et al., 2023). While anonymization tools offer some protection, LLMs can still detect subtle cues that reveal personal details.
Disclosure risks. LLMs’ human-like interactions foster user trust, increasing the likelihood of unintentional sharing of sensitive information (Kim and Sundar, 2012; Złotowski et al., 2015; Zhang et al., 2024). For example, Zhang et al. (Zhang et al., 2024) found that human-like interactions encourage users to unintentionally share sensitive and personally identifiable information with LLM-based conversational agents. They also noted that users believed they had to “pay the price” of privacy to get benefits from LLMs. This belief, compounded by manipulative interface designs (e.g., dark patterns (Lu et al., 2024)), amplifies disclosure risks.
These risks stem from users sharing sensitive data with LLMs directly or indirectly. We aim to develop a tool to inform users about privacy policies and potential risks before their data is shared.
@@ 2.2. Existing Methods for Protecting User Privacy when using LLM-enabled Applications
Existing methods to protect users’ privacy in LLM-enabled applications can be broadly categorized into model-side and user-side solutions. Model-side solutions operate at three stages (Smith et al., 2023): pre-processing, training, and post-training. Pre-processing involves data sanitization to remove sensitive information from training datasets, typically using automated techniques like pattern-based parsing (Lison et al., 2021; Vakili et al., 2022). However, defining private information is challenging due to its context-dependent nature and varied formats, making it difficult to guarantee complete sanitization (Brown et al., 2020). Training often employs differential privacy, which adds noise to data to prevent individual identification while retaining overall utility (Dwork et al., 2016). While effective in reducing memorization risks (Downey et al., 2022; Ponomareva et al., 2023), it can degrade model performance and increase computational demands (Ishihara, 2023). Post-training methods include filtering outputs with sensitive information (Perez et al., 2022) and knowledge unlearning, which forces models to forget specific data (Jang et al., 2022). However, these approaches face challenges in maintaining output diversity and ensuring consistent performance across use cases (Ishihara, 2023; Smith et al., 2023).
User-side solutions consider user engagement and interaction. Tools like ProPILE (Kim et al., 2024) let users test for personally identifiable information (PII) leaks in LLMs by probing their own data. However, these tools often lack contextual alignment with actual user scenarios. To address this, self-disclosure abstraction models (Dou et al., 2023) rephrase sensitive inputs into generalized terms while preserving utility (e.g., “I live in New Mexico” becomes “I live in the Southwest”). Although effective in reducing specificity, these methods do not fully inform users about potential privacy risks or data access.
In summary, existing methods such as privacy-preserving prompting (Edemacu and Wu, 2024) and automated consent mechanisms (Porcelli et al., 2023) focus on automating privacy tasks and minimizing user effort. We aim to promote user autonomy by supporting informed, context-aware decisions and enhancing privacy literacy. These approaches =complement each other, that is, existing automated tools can handle straightforward cases, while our approach aims to guide users through complex, ambiguous scenarios that require active intervention.
@@ 2.3. Contextual Privacy Policy
Contextual privacy policy (CPP) deconstructs lengthy privacy notices into shorter and context-specific ones that are relevant (Feth, 2017; Windl et al., 2022). Bergmann (Bergmann, 2008) showed that presenting only relevant information from privacy policies can significantly increase users’ privacy awareness. This also aligns with Nissenbaum’s theory of contextual integrity (Nissenbaum, 2004), which emphasizes that context is critical to determine whether a specific action is a violation of privacy.
CPP has been explored in web and mobile systems. In the general web domain, Ortlof et al. (Ortloff et al., 2020) developed a concept showcase of CPP for seven different websites and collected the design implications of CPP. Building on it, Windl et al. (Windl et al., 2022) presented PrivacyInjector, a more scalable system that can automatically generate and display CPPs. In the mobile domain, Pan et al. (Pan et al., 2024) proposed SeePrivacy, which integrates vision-based GUI understanding with privacy policy analysis to automatically generate CPPs for mobile applications.
Compared with previous work, our work addresses the unique challenges of implementing CPP for LLM-enabled applications. We not only create CPP for users’ sensitive input but also utilize LLMs to highlight potential privacy risks, helping users understand the implications of sharing sensitive information.
@@ 3. Co-Design Workshops
To understand user needs and explore the design space of tools to inform participants of privacy risks in LLM-enabled end-user applications, we conducted five co-design workshops [^1] The protocol of the workshop has been reviewed and approved by the IRB at our institution., including three in-person workshops and two online workshops.
@@ 3.1. Participants
“As noted above, we may use Content you provide us to improve our Services, for example, to train the models that power ChatGPT.”
@@ 3.2. Workshop Procedure
The workshop procedure is inspired by prior work in the privacy literature (Yao et al., 2019a, b; Lu et al., 2024). Each co-design workshop began with an introductory session, including ice-breaker activities to foster a collaborative atmosphere. Participants then engaged in a discussion on privacy by reading three news articles on incidents relevant to LLMs (e.g., an article entitled “ChatGPT maker OpenAI faces a lawsuit over how it used people’s data [^2] https://www.washingtonpost.com/technology/2023/06/28/openai-chatgpt-lawsuit-class-action/ from the Washington Post)”. They identified potential privacy risks and discussed strategies to mitigate them. This initial activity aimed to build a foundational understanding of privacy challenges associated with LLMs.
@@
Participants then engaged in an ideation session to share concerns and opportunities for privacy protection. This transitioned into three activities where they crafted privacy policy elements, assessed risks, and explored privacy incidents through sketching, writing, and discussions (Table 2).
@@
In the first activity, based on an example prompt input into the ChatGPT system, participants designed the content and presentation of retrieved privacy policy snippets that were relevant to sensitive information in example scenarios. Participants could either write descriptions of the interface features or create sketches of the interfaces. The purpose of this ideation activity was to gather user-specific requirements for contextual privacy policies in the context of LLM-enabled applications.
@@
The second activity expanded on the first one by asking participants to examine the prompts, privacy policy snippets, and associated privacy risks. The privacy risks were selected based on Lee et al.’s taxonomy of AI privacy risks (Lee et al., 2024) and were presented through detailed descriptions and visual aids, including examples of how personal information entered by users could be shared, used for advertising, or stored by service providers. Participants discussed and designed how to organize and present this information to better raise user awareness and support informed actions during user interaction with LLM-enabled applications. The activity specifically focused on the realistic context of use and how such information can fit into the user’s overall task that they are trying to achieve through the use of LLM-enabled applications. The goal was to understand whether highlighting LLM-related privacy risks could enhance users’ privacy awareness, and if so, how.
@@
The final activity focused on presenting specific privacy incidents, building on the risks identified earlier. Participants designed how real-world examples could highlight privacy implications. The objective was to understand the impact of explicit privacy incident examples on participants’ awareness and to determine effective methods for communicating these incidents within the context of LLM application use.